# Existence and non-existence of consistent estimators in supercritical controlled branching processes

### v1

@@ 5 Proofs
    $\mathbbm{E}\bigg{(}\frac{4}{n^{2}\cdot\varepsilon^{2}(Z_{n-1})}\cdot\big{(}\text{Var}(Z_{n}|\mathcal{F}_{n-1})\big{)}^{2}\bigg{)}=\mathbbm{E}\bigg{(}\frac{4}{n^{2}\cdot\varepsilon^{2}(Z_{n-1})}\cdot\big{(}\sigma^{2}\cdot\varepsilon(Z_{n-1})+m^{2}\cdot\nu^{2}(Z_{n-1})\big{)}^{2}\bigg{)}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0,$
    (22) is shown.
@@
    Since $\frac{(m-\tilde{m}_{n})^{2}}{n}\sum_{k=1}^{n}\varepsilon(Z_{k-1})$ is non-negative for all $n\in\mathbbm{N}_{1}$, we would like to show that
    $\mathbbm{E}\bigg{(}\frac{(m-\tilde{m}_{n})^{2}}{n}\sum_{k=1}^{n}\varepsilon(Z_{k-1})\bigg{)}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0,$ (23)
    because Markov’s inequality will then imply that $\frac{(m-\tilde{m}_{n})^{2}}{n}\sum_{k=1}^{n}\varepsilon(Z_{k-1})\stackrel{{\scriptstyle P}}{{\longrightarrow}}0$.
    We can write
    $\mathbbm{E}\bigg{(}\frac{(m-\tilde{m}_{n})^{2}}{n}\sum_{k=1}^{n}\varepsilon(Z_{k-1})\bigg{)}=\frac{1}{n}\cdot\mathbbm{E}\bigg{(}\frac{\text{Var}(Z_{n}|\mathcal{F}_{n-1})}{\varepsilon(Z_{n-1})}\cdot\sum_{k=1}^{n}\frac{\varepsilon(Z_{k-1})}{\varepsilon(Z_{n-1})}\bigg{)}.$ (24)
    Since $\{Z_{n},z_{0}\}$ is supercritical, for $t$ such that $1<t<\liminf_{z\to\infty}\tau(z)$ there exists $M\in\mathbbm{N}_{1}$ such that $\varepsilon(z)\cdot m>t\cdot z$ for all $z\geq M$. Given we are conditioned on the set of unbounded growth $\{\{Z_{n},z_{0}\}_{n}\to\infty\}$, there must therefore exist a finite $N$ such that $Z_{k}\geq M$ for all $k\geq N$. Then
    $\lim_{n\to\infty}\mathbbm{E}\bigg{(}\sum_{k=1}^{n}\frac{\varepsilon(Z_{k-1})}{\varepsilon(Z_{n-1})}\bigg{)}=\mathbbm{E}\bigg{(}\sum_{k=1}^{N-1}\frac{\varepsilon(Z_{k-1})}{\varepsilon(Z_{n-1})}\bigg{)}+\lim_{n\to\infty}\mathbbm{E}\bigg{(}\sum_{k=N}^{n}\frac{\varepsilon(Z_{k-1})}{\varepsilon(Z_{n-1})}\bigg{)},$
    where $\mathbbm{E}\Big{(}\sum_{k=1}^{N-1}\frac{\varepsilon(Z_{k-1})}{\varepsilon(Z_{n-1})}\Big{)}<\infty$ and since
    $\mathbbm{E}(Z_{n}|\mathcal{F}_{k-1})=\mathbbm{E}\big{(}\mathbbm{E}(Z_{n}|\mathcal{F}_{n-1})\big{|}\mathcal{F}_{k-1}\big{)}>mt\cdot\mathbbm{E}\big{(}Z_{n-1}\big{|}\mathcal{F}_{k-1}\big{)}>\dots>m\cdot t^{n-k}\cdot\varepsilon(Z_{k-1}),$
    we see
    $\lim_{n\to\infty}\mathbbm{E}\bigg{(}\sum_{k=N}^{n}\frac{\varepsilon(Z_{k-1})}{\varepsilon(Z_{n-1})}\bigg{)}$ $<\lim_{n\to\infty}\sum_{k=N}^{n}\frac{1}{m\cdot t^{n-k}}\cdot\mathbbm{E}\bigg{(}\frac{\mathbbm{E}(Z_{n}|\mathcal{F}_{k-1})}{\varepsilon(Z_{n-1})}\bigg{)}$
    $=\lim_{n\to\infty}\sum_{k=N}^{n}\frac{1}{m\cdot t^{n-k}}\cdot\mathbbm{E}\bigg{(}\mathbbm{E}\bigg{(}\frac{Z_{n}}{\varepsilon(Z_{n-1})}\bigg{|}\mathcal{F}_{n-1}\bigg{)}\bigg{)}$
    $=\lim_{n\to\infty}\sum_{k=N}^{n}\frac{1}{t^{n-k}}$
    $=\sum_{j=0}^{\infty}t^{-j}$
    $=\frac{t}{t-1},$
    so that $\lim_{n\to\infty}\mathbbm{E}\Big{(}\sum_{k=1}^{n}\frac{\varepsilon(Z_{k-1})}{\varepsilon(Z_{n-1})}\Big{)}$ is bounded from above by a constant. Consequently, since $\varepsilon(z)=\Theta(z)$ and $\nu^{2}(z)=O(z)$,
    $\lim_{n\to\infty}\mathbbm{E}\bigg{(}\frac{\text{Var}(Z_{n}|\mathcal{F}_{n-1})}{\varepsilon(Z_{n-1})}\cdot\sum_{k=1}^{n}\frac{\varepsilon(Z_{k-1})}{\varepsilon(Z_{n-1})}\bigg{)}=\lim_{n\to\infty}\mathbbm{E}\bigg{(}\frac{\sigma^{2}\cdot\varepsilon(Z_{n-1})+m^{2}\cdot\nu^{2}(Z_{n-1})}{\varepsilon(Z_{n-1})}\cdot\sum_{k=1}^{n}\frac{\varepsilon(Z_{k-1})}{\varepsilon(Z_{n-1})}\bigg{)}<\infty,$
    and thus, combining this with (24), the claim of (23) follows.
@@
    We want to show that $\frac{m^{2}-\tilde{m}^{2}_{n}}{n}\sum_{k=1}^{n}\frac{\nu^{2}(Z_{k-1})}{\varepsilon(Z_{k-1})}\stackrel{{\scriptstyle P}}{{\longrightarrow}}0$. Since $\varepsilon(z)=\Theta(z)$ and $\nu^{2}(z)=O(z)$, we know that $\frac{\nu^{2}(z)}{\varepsilon(z)}\leq C_{3}$ for all $z$, so, for any $n\in\mathbbm{N}_{1}$,
    $\bigg{|}\frac{m^{2}-\tilde{m}^{2}_{n}}{n}\sum_{k=1}^{n}\frac{\nu^{2}(Z_{k-1})}{\varepsilon(Z_{k-1})}\bigg{|}\leq C_{3}\cdot|m^{2}-\tilde{m}^{2}_{n}|.$
@@
    We have argued in (II) that there exist $M\in\mathbbm{N}_{1},\;t>1$ such that $\varepsilon(Z_{n-1})>\frac{tZ_{n-1}}{m}$ for all $n>M$. Then, given that $\nu^{2}(z)=O(z)$,
    $\lim_{n\to\infty}\text{Var}(\tilde{m}_{n})=\lim_{n\to\infty}\frac{m\sigma^{2}}{t}\cdot\mathbbm{E}(Z^{-1}_{n-1})+\lim_{n\to\infty}\frac{bm^{4}}{t^{2}}\cdot\mathbbm{E}(Z^{-1}_{n-1})=0,$
    and thus $\tilde{m}_{n}\stackrel{{\scriptstyle P}}{{\longrightarrow}}m$. It follows by the continuous mapping theorem that $m^{2}-\tilde{m}^{2}_{n}\stackrel{{\scriptstyle P}}{{\longrightarrow}}0$. Thus we see that
@@
Since we have shown that all three terms converge to zero in probability, we have shown that $\hat{\sigma}^{2}_{n}\stackrel{{\scriptstyle P}}{{\longrightarrow}}\sigma^{2}$. $\blacksquare$
5.3 Proofs for Section 3.3
Definition 5.4.
[…]

### v2

@@ 6 Proofs of consistency results
      $\mathbbm{E}\bigg{(}\frac{Z_{n}-m\varepsilon(Z_{n-1})}{n^{1/4}\cdot\sqrt{\varepsilon(Z_{n-1})}}\bigg{)}=0\quad\text{and}\quad\text{Var}\bigg{(}\frac{Z_{n}-m\varepsilon(Z_{n-1})}{n^{1/4}\cdot\sqrt{\varepsilon(Z_{n-1})}}\bigg{)}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0,$
      so, as a consequence of Chebyshev’s inequality, it follows that
      $\frac{Z_{n}-m\varepsilon(Z_{n-1})}{n^{1/4}\cdot\sqrt{\varepsilon(Z_{n-1})}}\stackrel{{\scriptstyle P}}{{\longrightarrow}}0.$
    3. (c)
      Defining
      $U_{n}:=Z_{n}-m\varepsilon(Z_{n-1})\quad\text{and}\quad J_{n}:=n^{3/4}\sqrt{\sum_{k=1}^{n}\varepsilon(Z_{k-1}}),$
      we can see that $\{U_{n}\}_{n\in\mathbbm{N}_{0}}$ is a martingale difference sequence, since $\mathbbm{E}(U_{n}|\mathcal{F}_{n-1})=0$ for all $n\in\mathbbm{N}_{1}$ and
      $\mathbbm{E}|U_{n}|$ $\leq\sqrt{\mathbbm{E}(U_{n}^{2})}$ (Hölder’s inequality)
      $=\sqrt{\mathbbm{E}(\text{Var}(Z_{n}|\mathcal{F}_{n-1}))}$
      $\leq\sqrt{(\sigma^{2}a+m^{2}b)\cdot\mathbbm{E}(Z_{n-1})}$ (by 6)
      $<\infty,$
      and $\{J_{n}\}_{n\in\mathbbm{N}_{0}}$ is positive, non-decreasing, with $\lim_{n\to\infty}J_{n}=\infty$. Additionally, using (6) and (27),
      $\sum_{n=1}^{\infty}J_{n}^{-2}\mathbbm{E}(U_{n}^{2}|\mathcal{F}_{n-1})$ $=\sum_{n=1}^{\infty}\frac{\text{Var}(Z_{n}|\mathcal{F}_{n-1})}{n^{3/2}\cdot\sum_{k=1}^{n}\varepsilon(Z_{k-1})}$
      $<\sum_{n=1}^{\infty}\frac{\text{Var}(Z_{n}|\mathcal{F}_{n-1})}{n^{3/2}\cdot\varepsilon(Z_{n-1})}$
      $\leq\sum_{n=1}^{\infty}\frac{\sigma^{2}+m^{2}C}{n^{3/2}}$
      $<\infty\;\;\text{a.s.},$
      so Theorem 6.1 applies, such that $J_{n}^{-1}\sum_{k=1}^{n}U_{k}=\frac{\sum_{k=1}^{n}(Z_{k}-m\varepsilon(Z_{k-1}))}{n^{3/4}\sqrt{\sum_{k=1}^{n}\varepsilon(Z_{k-1})}}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0$ as $n\to\infty$.
    Then, since $(a)$ is asymptotically bounded by a constant in probability, and $(b)$ and $(c)$ both converge to zero in probability, $\frac{2(m-\tilde{m}_{n})}{n}\sum_{k=1}^{n}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}\stackrel{{\scriptstyle P}}{{\longrightarrow}}0$.
@@
    We can decompose
    $\frac{(m-\tilde{m}_{n})^{2}}{n}\sum_{k=1}^{n}\varepsilon(Z_{k-1})=\frac{(Z_{n}-m\varepsilon(Z_{n-1}))^{2}}{n\cdot\varepsilon(Z_{n-1})}\cdot\sum_{k=1}^{n}\frac{\varepsilon(Z_{k-1})}{\varepsilon(Z_{n-1})},$
    where we know from Lemma 6.4 and (31) that, on $\{Z_{n}\to\infty\}$ and for s,ts,\,t such that $1<s<t<\liminf_{z\to\infty}\tau(z)$,
    $\lim_{n\to\infty}\mathbbm{P}\Bigg{(}\sum_{k=1}^{n}\frac{\varepsilon(Z_{k-1})}{\varepsilon(Z_{n-1})}\leq\frac{t}{am(1-s^{-1})}\Bigg{)}=1.$
    Therefore the result $\frac{(m-\tilde{m}_{n})^{2}}{n}\sum_{k=1}^{n}\varepsilon(Z_{k-1})\stackrel{{\scriptstyle P}}{{\longrightarrow}}0$ will follow if we show that
    $\frac{(Z_{n}-m\varepsilon(Z_{n-1}))^{2}}{n\cdot\varepsilon(Z_{n-1})}\stackrel{{\scriptstyle P}}{{\longrightarrow}}0.$ (32)
    Since $\frac{(Z_{n}-m\varepsilon(Z_{n-1}))^{2}}{n\cdot\varepsilon(Z_{n-1})}$ is non-negative, and given (27),
    $\mathbbm{E}\Bigg{(}\frac{(Z_{n}-m\varepsilon(Z_{n-1}))^{2}}{n\cdot\varepsilon(Z_{n-1})}\Bigg{)}=\mathbbm{E}\Bigg{(}\mathbbm{E}\bigg{(}\frac{(Z_{n}-m\varepsilon(Z_{n-1}))^{2}}{n\cdot\varepsilon(Z_{n-1})}\bigg{|}\mathcal{F}_{n-1}\bigg{)}\Bigg{)}\leq\frac{\sigma^{2}+m^{2}C}{n}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0.$
    Hence (32) follows from Markov’s inequality.
@@
    We want to show that $\frac{m^{2}-\tilde{m}^{2}_{n}}{n}\sum_{k=1}^{n}\frac{\nu^{2}(Z_{k-1})}{\varepsilon(Z_{k-1})}\stackrel{{\scriptstyle P}}{{\longrightarrow}}0$. Since there exists $C>0$ such that $\frac{\nu^{2}(z)}{\varepsilon(z)}\leq C$ for all $z$ by (27), for any $n\in\mathbbm{N}_{1}$,
    $\bigg{|}\frac{m^{2}-\tilde{m}^{2}_{n}}{n}\sum_{k=1}^{n}\frac{\nu^{2}(Z_{k-1})}{\varepsilon(Z_{k-1})}\bigg{|}\leq C\cdot|m^{2}-\tilde{m}^{2}_{n}|.$
@@
    We have argued in (I) that there exist $N\in\mathbbm{N}_{1},\;t>1$ such that $\varepsilon(Z_{n-1})>\frac{tZ_{n-1}}{m}$ for all $n>N$. Then, given that $\nu^{2}(z)\leq b\cdot z$ by (6),
    $\lim_{n\to\infty}\text{Var}(\tilde{m}_{n})\leq\lim_{n\to\infty}\frac{m\sigma^{2}}{t}\cdot\mathbbm{E}(Z^{-1}_{n-1})+\lim_{n\to\infty}\frac{bm^{4}}{t^{2}}\cdot\mathbbm{E}(Z^{-1}_{n-1})=0$
    on $\{Z_{n}\to\infty\}$, and thus $\tilde{m}_{n}\stackrel{{\scriptstyle P}}{{\longrightarrow}}m$. It follows by the continuous mapping theorem that $m^{2}-\tilde{m}^{2}_{n}\stackrel{{\scriptstyle P}}{{\longrightarrow}}0$. Thus we see that
@@
Since we have shown that (I), (II), and (III) all converge to zero in probability, we have shown that $\hat{\sigma}^{2}_{n}\stackrel{{\scriptstyle P}}{{\longrightarrow}}\sigma^{2}$.
The proof of Theorem 3.10 follows the same arguments as that of Theorem 3.5, albeit in a simpler setting.
Proof of Theorem 3.10:.
Let $\{Z_{n},z_{0}\}$ be a supercritical CBP with unknown control function $\phi(\cdot)$ having $\varepsilon(z)=\alpha z$ and $\nu^{2}(z)=\beta z$, and with $\sigma^{2}$ finite. Take $\mathcal{F}_{n}$ to be the $\sigma$-algebra generated by $(Z_{0},\dots,Z_{n})$. In this setting, the conditional variance of $Z_{n}$ given $\mathcal{F}_{n-1}$ has the following form:
$\text{Var}(Z_{n}|\mathcal{F}_{n-1})=\sigma^{2}\alpha+m^{2}\beta.$ (33)
Throughout the proof, we make the simplifying assumption that $\mathbbm{P}(Z_{1}>0\,|\,Z_{0}=z)=1$ for all $z\in\mathbbm{N}_{1}$, so that $I_{n}=\{1,\dots,n\}$. Under this assumption, $\mathbbm{P}(Z_{n}\to\infty)=1$, and the estimators simplify to
$\hat{g}_{n}:=\frac{1}{n}\sum_{k=1}^{n}\frac{Z_{k}}{Z_{k-1}},$
$\bar{h}_{n}:=\frac{1}{n}\sum_{k=1}^{n}\frac{(Z_{k}-m\alpha\cdot Z_{k-1})^{2}}{Z_{k-1}},\quad\text{and}\quad\hat{h}_{n}:=\frac{1}{n}\sum_{k=1}^{n}\frac{(Z_{k}-\tilde{g}_{n}\cdot Z_{k-1})^{2}}{Z_{k-1}}.$
The extension of the proof to the case where there exists $z\in\mathbbm{N}_{1}$ such that $\mathbbm{P}(Z_{1}=0\,|\,Z_{0}=z)>0$ follows the same arguments as in the proof of Theorem 3.5, part (i), case (b). To avoid repeating ourselves, we omit it.
(i) Strong consistency of $\boldsymbol{\hat{g}_{n}}$: Let $U_{n}:=\tilde{g}_{n}-m\alpha$, where we recall that $\tilde{g}_{n}:=Z_{n}/Z_{n-1}$. We then have $\mathbbm{E}(U_{n})=\mathbbm{E}(\mathbbm{E}(U_{n}|\mathcal{F}_{n-1}))=0$ and $\mathbbm{E}|U_{n}|\leq 2m\alpha$, while we can use (33) to see that
$\mathbbm{E}(U_{n}^{2}|\mathcal{F}_{n-1})=Z_{n-1}^{-2}\cdot\text{Var}(Z_{n}|\mathcal{F}_{n-1})=\frac{\sigma^{2}\alpha+m^{2}\beta}{Z_{n-1}}.$
Our assumption that $\mathbbm{P}(Z_{1}>0|Z_{0}=z)=1$ for all $z\in\mathbbm{N}_{1}$ implies that, for all $k\in\mathbbm{N}_{1}$, $\mathbbm{P}(Z_{k}=0|Z_{0}=z)=0$, and thus that $Z_{k}\geq 1$ a.s.. Hence
$\frac{\sigma^{2}\alpha+m^{2}\beta}{Z_{n-1}}\leq\sigma^{2}\alpha+m^{2}\beta,$
and therefore
$\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}(U_{n}^{2}|\mathcal{F}_{n-1})\leq\sum_{n=1}^{\infty}\frac{\sigma^{2}\alpha+m^{2}\beta}{n^{2}}<\infty.$
Then, by Corollary 6.2, $\frac{1}{n}\sum_{k=1}^{n}U_{k}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0$ and thus $\hat{g}_{n}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}m\alpha$.
(ii) Strong consistency of $\boldsymbol{\bar{h}_{n}}$: Under the assumption that $m\alpha$ is known, we introduce the random variable
$U_{n}:=\frac{(Z_{n}-m\alpha\cdot Z_{n-1})^{2}}{Z_{n-1}}-\sigma^{2}\alpha-m^{2}\beta,$
for which we see that
$\mathbbm{E}(U_{n}|\mathcal{F}_{n-1})=Z_{n-1}^{-1}\cdot\text{Var}(Z_{n}|\mathcal{F}_{n-1})-\sigma^{2}\alpha-m^{2}\beta=0,$
so that consequently $\mathbbm{E}(U_{n})=0$, and
$\mathbbm{E}|U_{n}|\leq\mathbbm{E}\Bigg{|}\frac{(Z_{n}-m\alpha\cdot Z_{n-1})^{2}}{Z_{n-1}}\Bigg{|}+\sigma^{2}\alpha+m^{2}\beta=2\sigma^{2}\alpha+2m^{2}\beta<\infty.$
In addition, by a lengthy yet elementary expansion, we can find that
$\mathbbm{E}(U_{n}^{2}|\mathcal{F}_{n-1})$ $=\frac{m^{4}\mathbbm{E}(\phi(Z_{n-1})-\varepsilon(Z_{n-1}))^{4}}{Z_{n-1}^{2}}+\frac{6\sigma^{2}m^{2}\iota(Z_{n-1})}{Z_{n-1}^{2}}+\frac{\alpha\mathbbm{E}(\xi-m)^{4}}{Z_{n-1}}$
$\quad+\frac{4\gamma m\beta+3\sigma^{4}\beta+18\sigma^{2}m^{2}\beta+m^{4}\alpha^{4}-3\sigma^{4}\alpha}{Z_{n-1}}+3\sigma^{4}\alpha^{2}-12\sigma^{2}m^{2}\alpha\beta+12\sigma^{2}m^{2}\alpha^{2}$
$\quad+4m^{4}\alpha^{2}-(\sigma^{2}\alpha+m^{2}\beta)^{2}.$
Given our assumptions on the bounds on $\mathbbm{E}(\phi(Z_{n-1})-\varepsilon(Z_{n-1}))^{4}$ and $\mathbbm{E}(\xi-m)^{4}$, there will therefore exist a positive constant $C$ such that $\mathbbm{E}\big{(}U_{n}^{2}\big{|}\mathcal{F}_{n-1}\big{)}\leq C$. Hence
$\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}(U_{n}^{2}|\mathcal{F}_{n-1})\leq\sum_{n=1}^{\infty}\frac{C}{n^{2}}<\infty.$
Then by Corollary 6.2, $\frac{1}{n}\sum_{k=1}^{n}U_{k}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0$ and thus $\bar{h}_{n}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\sigma^{2}\alpha+m^{2}\beta$.
(iii) Weak consistency of $\boldsymbol{\hat{h}_{n}}$: We abandon the assumption that $m\alpha$ is known. We can decompose
$\hat{h}_{n}=\bar{h}_{n}+\underbrace{\frac{2(m\alpha-\tilde{g}_{n})}{n}\sum_{k=1}^{n}(Z_{k}-m\alpha Z_{k-1})}_{\text{I}}+\underbrace{\frac{(m\alpha-\tilde{g}_{n})^{2}}{n}\sum_{k=1}^{n}Z_{k-1},}_{\text{II}}$
where we have shown in (ii) that $\bar{h}_{n}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\sigma^{2}\alpha+m^{2}\beta$. It remains to show that I and II both converge in probability to zero.
  1. (I)
    We can decompose
    $\frac{2(m\alpha-\tilde{g}_{n})}{n}\sum_{k=1}^{n}(Z_{k}-m\alpha Z_{k-1})=-2\underbrace{\sqrt{\frac{\sum_{k=1}^{n}Z_{k-1}}{Z_{n-1}}}}_{\text{(a)}}\cdot\underbrace{\frac{Z_{n}-m\alpha Z_{n-1}}{n^{1/4}\cdot\sqrt{Z_{n-1}}}}_{\text{(b)}}\cdot\underbrace{\frac{\sum_{k=1}^{n}(Z_{k}-m\alpha Z_{k-1})}{n^{3/4}\sqrt{\sum_{k=1}^{n}Z_{k-1}}}}_{\text{(c)}},$
    and consider each of the components $(a)$, $(b)$ and $(c)$ in turn.
    1. (a)
      As a result of Lemma 6.4, for any $s$ such that $1<s<m\alpha$, we have that
      $\lim_{n\to\infty}\mathbbm{P}\Bigg{(}\sqrt{\frac{\sum_{k=1}^{n}Z_{k-1}}{Z_{n-1}}}\leq\sqrt{\frac{1}{1-s^{-1}}}\Bigg{)}=1,$
      that is, $\sqrt{\frac{\sum_{k=1}^{n}Z_{k-1}}{Z_{n-1}}}$ is asymptotically bounded by a constant in probability.
    2. (b)
      Given that
      $\mathbbm{E}\bigg{(}\frac{Z_{n}-m\alpha Z_{n-1}}{n^{1/4}\cdot\sqrt{Z_{n-1}}}\bigg{|}\mathcal{F}_{n-1}\bigg{)}=0\quad\text{and}\quad\text{Var}\bigg{(}\frac{Z_{n}-m\alpha Z_{n-1}}{n^{1/4}\cdot\sqrt{Z_{n-1}}}\bigg{|}\mathcal{F}_{n-1}\bigg{)}=\frac{\sigma^{2}\alpha+m^{2}\beta}{\sqrt{n}},$
      we see that
      $\mathbbm{E}\bigg{(}\frac{Z_{n}-m\alpha Z_{n-1}}{n^{1/4}\cdot\sqrt{Z_{n-1}}}\bigg{)}=0\quad\text{and}\quad\text{Var}\bigg{(}\frac{Z_{n}-m\alpha Z_{n-1}}{n^{1/4}\cdot\sqrt{Z_{n-1}}}\bigg{)}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0,$
      so, by Chebyshev’s inequality, it follows that
      $\frac{Z_{n}-m\alpha Z_{n-1}}{n^{1/4}\cdot\sqrt{Z_{n-1}}}\stackrel{{\scriptstyle P}}{{\longrightarrow}}0.$
    3. (c)
      Defining
      $U_{n}:=Z_{n}-m\alpha Z_{n-1}\quad\text{and}\quad J_{n}:=n^{3/4}\sqrt{\sum_{k=1}^{n}Z_{k-1}},$
      we can see that $\{U_{n}\}_{n\in\mathbbm{N}_{0}}$ is a martingale difference sequence, since $\mathbbm{E}(U_{n}|\mathcal{F}_{n-1})=0$ for all $n\in\mathbbm{N}_{1}$ and
      $\mathbbm{E}|U_{n}|$ $\leq\sqrt{\mathbbm{E}(U_{n}^{2})}$ (Hölder’s inequality)
      $=\sqrt{\mathbbm{E}(\text{Var}(Z_{n}|\mathcal{F}_{n-1}))}$
      $=\sqrt{(\sigma^{2}\alpha+m^{2}\beta)\cdot\mathbbm{E}(Z_{n-1})}$
      $<\infty,$
      and $\{J_{n}\}_{n\in\mathbbm{N}_{0}}$ is positive, non-decreasing, with $\lim_{n\to\infty}J_{n}=\infty$. Additionally,
      $\sum_{n=1}^{\infty}J_{n}^{-2}\mathbbm{E}(U_{n}^{2}|\mathcal{F}_{n-1})$ $=\sum_{n=1}^{\infty}\frac{\text{Var}(Z_{n}|\mathcal{F}_{n-1})}{n^{3/2}\cdot\sum_{k=1}^{n}Z_{k-1}}$
      $=\sum_{n=1}^{\infty}\frac{(\sigma^{2}\alpha+m^{2}\beta)\cdot Z_{n-1}}{n^{3/2}\cdot\sum_{k=1}^{n}Z_{k-1}}$
      $<\sum_{n=1}^{\infty}\frac{(\sigma^{2}\alpha+m^{2}\beta)}{n^{3/2}}$
      $<\infty\;\;\text{a.s.},$
      so Theorem 6.1 applies, such that $J_{n}^{-1}\sum_{k=1}^{n}U_{k}=\frac{\sum_{k=1}^{n}(Z_{k}-m\alpha Z_{k-1})}{n^{3/4}\sqrt{\sum_{k=1}^{n}Z_{k-1}}}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0$ as $n\to\infty$.
    Then, since $(a)$ is asymptotically bounded by a constant in probability, and $(b)$ and $(c)$ both converge to zero in probability, $\frac{2(m\alpha-\tilde{g}_{n})}{n}\sum_{k=1}^{n}(Z_{k}-m\alpha Z_{k-1})\stackrel{{\scriptstyle P}}{{\longrightarrow}}0$.
  2. (II)
    We can decompose
    $\frac{(m\alpha-\tilde{g}_{n})^{2}}{n}\sum_{k=1}^{n}Z_{k-1}=\frac{(Z_{n}-m\alpha Z_{n-1})^{2}}{n\cdot Z_{n-1}}\cdot\sum_{k=1}^{n}\frac{Z_{k-1}}{Z_{n-1}},$
    and have already shown in Lemma 6.4 that $\sum_{k=1}^{n}\frac{Z_{k-1}}{Z_{n-1}}$ is asymptotically bounded in probability by a constant. Therefore the result $\frac{(m\alpha-\tilde{g}_{n})^{2}}{n}\sum_{k=1}^{n}Z_{k-1}\stackrel{{\scriptstyle P}}{{\longrightarrow}}0$ will follow if we show that $\frac{(Z_{n}-m\alpha Z_{n-1})^{2}}{n\cdot Z_{n-1}}\stackrel{{\scriptstyle P}}{{\longrightarrow}}0$. Noting that $\frac{(Z_{n}-m\alpha Z_{n-1})^{2}}{n\cdot Z_{n-1}}$ is non-negative, and that
    $\mathbbm{E}\Bigg{(}\frac{(Z_{n}-m\alpha Z_{n-1})^{2}}{n\cdot Z_{n-1}}\Bigg{)}=\mathbbm{E}\Bigg{(}\mathbbm{E}\bigg{(}\frac{(Z_{n}-m\alpha Z_{n-1})^{2}}{n\cdot Z_{n-1}}\bigg{|}\mathcal{F}_{n-1}\bigg{)}\Bigg{)}=\frac{\sigma^{2}\alpha+m^{2}\beta}{n}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0,$
    this follows from Markov’s inequality.
Theorem 3.13 will require the use of the following lemma, which we state without proof, noting that its proof is effectively unchanged from that of Lemma 6.4, replacing $Z_{n}$ with $\phi(Z_{n})$. To help illustrate this, note that, given that $\varepsilon(z)=\alpha z$ for all $z\in\mathbbm{N}_{0}$,
$\mathbbm{E}(Z_{n}\,|\,Z_{n-1})=m\alpha\cdot Z_{n-1}\quad\text{and}\quad\mathbbm{E}(\phi(Z_{n})\,|\,\phi(Z_{n-1}))=m\alpha\cdot\phi(Z_{n-1}).$